# Compactor: Calibrated Query-Agnostic KV Cache Compression with Approximate Leverage Scores

### v1

@@ 3.3 Attention-Based Scoring
where ${1}^{\top}\in\mathbb{R}^{1\times n}$ is a vector of all ones and $\mathbf{M}\in\mathbb{R}^{n}$ is an upper triangular matrix filled with $\infty$- ∞ to enforce causality, and selects the top-$k$ highest scoring tokens for retention. Similarly, SnapKV uses the attention of the last $W$ queries (on all keys) to generate importance scores:
@@
where $\mathbf{Q}[i]$ (and similarly $\mathbf{K}[i]$) is the submatrix containing rows $(i-1)C+1$ through $iC$. We apply mean-pooling to $\vec{{a}}$ as suggested by [24] and value norm scaling [17] to all mechanisms tested).
@@ 3.4 Score Blending and KV Selection
The Compactor eviction procedure is then: (1) compute token scores $\vec{s}$ as described by Eq 6; (2) retain the top-$\lceil r\cdot N\rceil$ tokens with the highest scores. Note the scoring and eviction procedure is applied independently to each head and layer.
@@ 3.5 Context-Calibrated Compression
Let $\widetilde{c}_{r}$ be the same context after applying a compression technique with retention $r$. Let $q$ denote any question. Our goal is to predict, for any $(q,c,r)$ pair, the expected NLL ratio:
@@
for any text $t\in\mathcal{V}^{L}$. To make this possible in the query-agnostic setting, we make the following crucial assumption:
@@
for some function $f$, parameterized only by $c,r$; i.e., we assume the gain in NLL of any text $t$ conditioned on $\tilde{c}_{r}$ can be explained entirely by the original context and $r$. Such an assumption is necessary in the query-agnostic regime. Empirically (see Fig. 6) we find $g(r,c)$ decays smoothly and monotonically with $r$. As such, to approximate $g(r,q,c)\approx f(r,c)$, we fit the two-parameter curve
$k=\alpha\cdot\text{NLL}(c)+\beta\qquad f_{\alpha,\beta}(r,c)=\frac{\exp\bigl{(}rk-k\bigr{)}-\exp(-k)}{1-\exp(-k)}$
@@
where $q$ is the question being asked, $t$ is the ground truth answer, and $y$ is the ratio of the NLL of the true answer under the original and compressed contexts. At inference time, given a user-specified quality budget $\tau\in(0,1]$ and context $c$, (for example $\tau=0.95$ allows an NLL increase of at most $\frac{1}{0.95}\approx 5\%$), we choose $r^{\star}(c,\tau)=\arg\max_{r\in(0,1)}f_{\alpha,\beta}(r,c)$ such that $f_{\alpha,\beta}(r,c)\geq\tau$, for which there is a closed form solution. Compactor then keeps the top $\lceil r^{\star}\!\cdot N\rceil$ tokens per head. In practice, we use asymmetrical MSE to penalize under-estimation of $r^{\star}(c,\tau)$. In summary, we fit a curve to predict the level of degradation that will occur when retaining $r$ fraction of tokens under a particular compression method; at inference time, we use the inverse of the curve to determine what retention rate the given context can support without excessively degrading performance.
@@ 4 Experiments
RULER [18] is a synthetic benchmark that asks: given a stated context window, how much of it can the model actually use? It consists of 13 tasks organized into four categories: (1) needle-in-a-haystack (NIAH) retrieval with varying numbers of and types of “needles”, (2) multi-hop tracing that forces models to follow chains of references scattered through the text, (3) aggregation tasks that require fusing information from many locations, and (4) stress tests that add distractor tokens to short-form question tasks. We use the RULER-4k variant because (1) tested models demonstrate excellent performance with no compression, (2) models have poor performance at higher compression rates, and (3) we are chiefly concerned with performance as a function of compression rate, instead of number of tokens retained.
Longbench [3] is a benchmark designed to evaluate long-context understanding in LLMs. It consists of 21 datasets, covering the following task categories: single-document QA, multi-document QA, summarization, few-shot learning, synthetic tasks, and code completion. LongBench requires comprehension of extensive passages, and probes models’ ability to maintain coherence, retrieve relevant information, and reason over non-synthetic long inputs. We evaluate on the English subset, and provide results for tasks that have a well-defined query (which excludes code-completion; for completeness we include results on the coding portion of Longbench in Appendix  E)
Additionally, we assess the performance of context-calibrated compression on Longbench under two settings: Zero-Shot, and Finetuned. In the Zero Shot setting, for each document in the Longbench corpus, we predict the maximum compression rate the (eviction method, document) pair supports (see Sec  5.4). In the Finetuned setting, we finetune the base model on documents in the Longbench corpus (not queries or answers), and apply the aforementioned procedure to the finetuned model on the same Longbench corpus. Such an experiment is useful for assessing the compressive capability of the model when targeted for a particular corpus known a priori. The finetuning is performed with rank 128 LoRA adaptors on Q,K,V matrices for 8 epochs [19].
@@
Figure 3: Median wall-clock overhead of the selection mechanism for a single layer of Llama 3.1 (batch size 1).
We use the following parameters for all experiments: $\lambda=0.3$, sketching dimension $k=64$, chunk size $C=256$, $\tau=0.95$ for Llama 3.1 and $\tau=0.90$ for Qwen 2.5; $\lambda,C$ were chosen to maximize performance on a validation split of RULER when searching over $\lambda\in\{0.15,0.2,0.3,0.4,0.5,0.6\}$ and $C\in\{64,128,256,512\}$. We demonstrate the stability of Compactor across parameter choices in § 5.5. While the choice of $k=64$ violates the bounds required for Theorem 2, we empirically find the approximation is sufficient to achieve good performance. Such a phenomenon is not a new observation, and much prior work has shown that the bounds tend to be loose on real-world datasets [5, 30, 21]. All experiments are conducted in a query-agnostic manner.
@@ 5 Results
[…]
Figure 4: Performance of eviction methods on RULER NIAH and QA subtasks at 50% KV retention cache on Llama 3.1. Performance is measured as a fraction of full KV performance.
@@ 5.1 Selection Mechanism Speed
We benchmark the speed of different eviction mechanisms across various context lengths. Figure 3 shows wall-clock execution time of the eviction mechanism for a single layer for Llama 3.1 (with batch size 1). We plot the execution time of Flash Attention 2 to provide context as a lower bound on the amount of time spent in a layer. We see that Compactor is about as fast as SnapKV for context lengths above $16$k tokens; the overhead at short contexts is due to the time spent computing an SVD, but this time is constant for all context lengths.
We additionally show the cost of computing only outlier scores (Compactor (no $\vec{a}$)) to demonstrate the minimal overhead of approximate leverage score computation; the increase in execution time for Compactor (no $\vec{a}$) is almost entirely due to matrix multiplication (line 5 in Algorithm 1). As such, Compactor can be used as a drop-in replacement over other eviction mechanisms in the query-agnostic regime, with no performance overhead for long contexts.
@@ 5.2 RULER
Figure 5 plots mean RULER score (across all tasks) versus KV retention for Compactor and four baselines (SnapKV, $\textsf{H}_{2}\textsf{O}$, PyramidKV, random eviction) for both Llama 3.1 and Qwen 2.5. $100\%$ KV retention denotes the score of the model with no compression and is shown as a dashed line. We observe remarkably similar performance characteristics across both models. As the cache is pruned, Compactor degrades most gracefully: at $75\%$ retention it recovers $93.8\%$ of baseline, at $50\%$ it maintains $87.6\%$, and even at only $10\%$ retention it still achieves $59.5\%$. Performance when using SnapKV, PyramidKV, or $\textsf{H}_{2}\textsf{O}$ degrades linearly with KV retention setting, confirming the inadequacy of these methods in the query-agnostic regime. The performance of all methods is poor at severe compression ratios, demonstrating the inherent difficulty of the task. Performance of eviction methods on NIAH and QA tasks is shown in Figure 4, and demonstrates the variable performance that different tasks support.
[…]
@@
5.3 Longbench
@@
Single-Doc QA Compactor 0.484 0.485 0.434 0.279 0.172 0.438 0.433 0.369 0.277 0.174
SnapKV 0.407 0.290 0.200 0.136 0.356 0.234 0.184 0.140
PyramidKV 0.371 0.278 0.186 0.139 0.361 0.230 0.184 0.139
Multi-Doc QA Compactor 0.476 0.465 0.448 0.389 0.254 0.563 0.572 0.549 0.455 0.419
SnapKV 0.407 0.290 0.200 0.136 0.494 0.423 0.316 0.259
PyramidKV 0.371 0.278 0.186 0.139 0.490 0.411 0.297 0.264
Summarization Compactor 0.286 0.283 0.265 0.238 0.206 0.258 0.253 0.238 0.214 0.196
SnapKV 0.265 0.242 0.198 0.168 0.240 0.219 0.182 0.154
PyramidKV 0.264 0.239 0.197 0.172 0.242 0.212 0.176 0.140
Few Shot Compactor 0.490 0.509 0.492 0.400 0.364 0.617 0.623 0.588 0.497 0.377
SnapKV 0.493 0.441 0.327 0.290 0.593 0.551 0.482 0.342
PyramidKV 0.500 0.435 0.321 0.291 0.588 0.545 0.480 0.336
Synthetic Compactor 0.580 0.594 0.520 0.153 0.081 0.607 0.593 0.447 0.153 0.073
SnapKV 0.527 0.382 0.047 0.007 0.593 0.383 0.127 0.073
PyramidKV 0.533 0.380 0.094 0.013 0.590 0.377 0.13 0.069
Total Compactor 0.455 0.458 0.425 0.302 0.225 0.489 0.485 0.438 0.331 0.261
SnapKV 0.417 0.344 0.196 0.153 0.445 0.360 0.267 0.202
PyramidKV 0.413 0.337 0.201 0.146 0.444 0.353 0.264 0.198
We evaluate Compactor on Longbench to assess performance on real-world long context tasks. Table 1 reports per-task and average scores on LongBench for Compactor , SnapKV, and PyramidKV at $50\%$, $25\%$, $10\%$, and $5\%$ retention. $\textsf{H}_{2}\textsf{O}$ was excluded due to insufficient memory on long contexts. At $50\%$ retention, Compactor matches or slightly exceeds the full-cache performance, with particular gains in few-shot and synthetic tasks with less contextual diversity. Even at $25\%$, Compactor retains $93\%$ of the full-cache score, whereas SnapKV and PyramidKV retain $74\%$ of the baseline. Under more aggressive pruning ( $10\%$ and $5\%$), Compactor ’s advantage grows: it outperforms the next best baseline by over $50\%$ relative at $10\%$, and by over $40\%$ at $5\%$. These results confirm that Compactor reliably preserves tokens that drive task performance.
@@ 5.4 Context-Calibrated Compression
[…]
Figure 6: NLL vs KV retention on 13 tasks from RULER, one line per sub-task. For each sub task we compute the (reciprocal) gain in NLL of the ground-truth answer when conditioned on compressed contexts (for each compression method). This empirical trend motivates the exponential calibration function introduced in §3.5
0020202020404040406060606080808080100100100100FinetunedZero ShotFull (100%)Full (100%)Compactor  (38%)SnapKV (72%)PyramidKV (75%)Random (90%)Compactor (22%)SnapKV (44%)PyramidKV (51%)Random (76%)KV Retention (%)
Figure 7: The top bar shows the KV retention rates that each compression method induces when using context-calibrated compression on Longbench. The bottom bar shows the same when the LLM is finetuned on documents from the Longbench test set (no queries). In all cases, the performance of the compression methods is statistically indistinguishable from full KV cache performance. Results are shown for Llama-3.1 8B.
@@
For all tested methods, the performance when applying context-calibrated compression is indistinguishable from full (100% KV) cache performance (one way ANOVA, $p=0.53$ when performing 20 trials over the randomness in the mechanism). This confirms both the utility of NLL as a proxy for downstream task performance and the efficacy of the proposed calibration function at ensuring compression does not degrade performance. Note Compactor is able to achieve this performance while retaining significantly fewer tokens ($\approx 37\%$, while SnapKV retains $\approx 72\%$). Both SnapKV and PyramidKV provide only a marginal improvement over the random eviction baseline, far removed from the excellent performance they demonstrate in the query-aware setting [6, 24], where they can match baseline performance while retaining only $10\%$ of the cache.
Lastly, we finetune the model on documents from Longbench (no queries or answers), and evaluate the performance of context-calibrated compression with the finetuned model on Longbench (we evaluate and finetune with the same set of documents). We would expect that the NLL of the documents to fall, allowing for higher compression rates for the documents. Such a procedure is useful when we know a priori which corpus will be assessed, and we desire the highest possible accuracy and compression on contexts from this corpus. For example, a question-answering LLM may be finetuned on a corpus of supporting documents, while simultaneously being presented with relevant retrieved documents from that corpus at test time. The result of this experiment is shown in Fig 7 (“Finetuned"). As expected, we see that all models achieve higher compression rates when finetuned on the test corpus. Compactor still significantly outperforms competing techniques, though the gap is shortened; notably, Compactor without training (Zero Shot) outperforms SnapKV and PyramidKV even after finetuning

### v2

@@ 3.3 Attention-Based Scoring
where ${1}^{\top}\in\mathbb{R}^{1\times n}$ is a vector of all ones and $\mathbf{M}\in\mathbb{R}^{n}$ is an upper triangular matrix filled with -$\infty$ to enforce causality, and selects the top-$k$ highest scoring tokens for retention. Similarly, SnapKV uses the attention of the last $W$ queries (on all keys) to generate importance scores:
@@
where $\mathbf{Q}[i]$ (and similarly $\mathbf{K}[i]$) is the submatrix containing rows (i1)C+1(i-1)C+1 through $iC$. We apply mean-pooling to $\vec{{a}}$ as suggested by (li_snapkv_2024).
@@ 3.4 Score Blending and KV Selection
The Compactor eviction procedure is then: (1) compute token scores $\vec{s}$ as described by Eq 6; (2) retain the top-$\lceil r\cdot N\rceil$ tokens with the highest scores (either across all heads, or within a single a head). Note the scoring procedure is applied independently to each head and layer.
@@ 3.5 Context-Calibrated Compression
Let $\widetilde{c}_{r}$ be the same context after applying a compression technique with retention $r$. Let $q$ denote any question. Our goal is to predict, for any $(q,c,r)$ triple, the expected NLL ratio:
@@
for any text $t\in\mathcal{V}^{L}$. To make this possible in the query-agnostic setting we make the following crucial assumption:
@@
for some function $f$, parameterized only by $c,r$. I.e., we assume the gain in NLL of any text $t$ conditioned on $\tilde{c}_{r}$ can be explained entirely by the original context and $r$. Such an assumption is necessary in the query-agnostic regime. Empirically (see Fig. 5) we find $g(r,c)$ decays smoothly and monotonically with $r$. As such, to approximate $g(r,q,c)\approx f(r,c)$, we fit the two-parameter curve
$b=\alpha\cdot\text{NLL}(c)+\beta\qquad f_{\alpha,\beta}(r,c)=\frac{\exp\bigl(rb-b\bigr)-\exp(-b)}{1-\exp(-b)}$
@@
where $q$ is the question being asked, $t$ is the ground truth answer, and $y$ is the ratio of the NLL of the true answer under the original and compressed contexts. At inference time, given a user-specified quality budget $\tau\in(0,1]$ and context $c$, (for example $\tau=0.95$ allows an NLL increase of at most $\frac{1}{0.95}\approx 5\%$), we choose $r^{\star}(c,\tau)=\arg\max_{r\in(0,1)}f_{\alpha,\beta}(r,c)$ such that $f_{\alpha,\beta}(r,c)\geq\tau$, for which there is a closed form solution. Compactor then keeps the top $\lceil r^{\star}\!\cdot N\rceil$ tokens per head. In practice, we use asymmetrical MSE to penalize under-estimation of $r^{\star}(c,\tau)$ (i.e weight such errors higher). In summary, we fit a curve to predict the level of degradation that will occur when retaining $r$ fraction of tokens under a particular compression method. At inference time we use the inverse of the curve to determine what retention rate the given context can support without excessively degrading performance.
@@ 4 Experiments
RULER (hsieh_ruler_2024) is a synthetic benchmark that asks: given a stated context window, how much of it can the model actually use? It consists of 13 tasks organized into four categories: (1) needle‑in‑a‑haystack (NIAH) retrieval with varying numbers of and types of “needles”, (2) multi‑hop tracing that forces models to follow chains of references scattered through the text, (3) aggregation tasks that require fusing information from many locations, and (4) stress tests that add distractor tokens to short-form question tasks. We use the RULER-4k variant because (i) tested models demonstrate excellent performance with no compression, (ii) models have poor performance at higher compression rates, and (iii) we are chiefly concerned with performance as a function of compression rate, instead of number of tokens retained.
Longbench (bai_longbench_2024) is a benchmark designed to evaluate long‐context understanding in LLMs. It consists of 21 datasets, covering the following task categories: single‐document QA, multi‐document QA, summarization, few‐shot learning, synthetic tasks, and code completion. LongBench requires comprehension of extensive passages, and probes models’ ability to maintain coherence, retrieve relevant information, and reason over non-synthetic long inputs. We evaluate on the English subset.
Additionally, we assess the performance of context-calibrated compression on Longbench under two settings: Zero-Shot, and Finetuned. In the Zero Shot setting, for each document in the Longbench corpus, we predict the maximum compression rate the (eviction method, document) pair supports (see Sec  5.4). In the Finetuned setting, we finetune the base model on documents in the Longbench corpus (not including queries or answers), and apply the aforementioned procedure to the finetuned model on the same Longbench corpus. Such an experiment is useful for assessing the compressive capability of the model when targeted for a particular corpus known a priori. The finetuning is performed with rank 128 LoRA adaptors on Q,K,V matrices for 3 epochs (hu_lora_2022).
@@
Figure 2: Median wall-clock overhead of the selection mechanism for one layer of Llama 3.1 (batch size 1).
TOVAKeyDiffSnapKVCompactor 020%40%60%80%100%Performance (% of Full KV)NIAHQA
Figure 3: Performance of eviction methods on RULER NIAH UUID and QA subtasks at 50% KV retention cache on Llama 3.1.
We use the following parameters for all experiments: $\lambda=0.3$, sketching dimension $k=48$, chunk size $C=256$, $\tau=0.95$ for Llama 3.1 and $\tau=0.90$ for Qwen 2.5. We demonstrate the stability of Compactor across parameter choices in § 5.5. While the choice of $k=48$ violates the bounds required for Theorem 2, we empirically find the approximation is sufficient to achieve good performance. Such a phenomenon is not a new observation, and much prior work has shown that the bounds tend to be loose on real-world datasets (broadbent_subset_2010; paschou_pca-correlated_2007; jolliffe_discarding_1972). All experiments are conducted in a query-agnostic manner, and with head-adaptive compression (feng_ada-kv_2025) (except 5, which is computed without head-adaptive compression to better qualitatively demonstrate task-variability). We compare Compactor to TOVA (oren_transformers_2024), KeyDiff (park_keydiff_2025), SnapKV (li_snapkv_2024), DuoAttention (xiao_duoattention_2024), and PyramidKV (cai_pyramidkv_2024) as baselines. We exclude $\textsf{H}_{2}\textsf{O}$ given its computational expense (see Figure 3). We use greedy sampling for all results.
@@ 5.1 Selection Mechanism Speed
We benchmark the speed of different eviction mechanisms across various context lengths. Figure 3 shows wall-clock execution time of the eviction mechanism for a single layer for Llama 3.1 (with batch size 1). We plot the execution time of Flash Attention 2 to provide context as a lower bound on the amount of time spent in a layer. We see that Compactor is about as fast as SnapKV for context lengths above 1616k tokens; the overhead at short contexts is due to the time spent computing an SVD, but this time is constant for all context lengths.
We additionally show the cost of computing only outlier scores (Compactor ($\vec{o}$)) to demonstrate the minimal overhead of approximate leverage score computation; the increase in execution time for Compactor (no $\vec{a}$) is almost entirely due to matrix multiplication (line 5 in Algorithm 1). As such, Compactor can be used as a drop-in replacement over other eviction mechanisms in the query-agnostic regime, with very little performance overhead for long contexts.
@@ 5.2 RULER
0202040406060808010010005050100100KV Retention ($\%$)Mean RULER ScoreLlama 3.1BaseCompactorSnapKVTOVAPyramidKVRandomDuoAttention020204040606080801001002020404060608080100100KV Retention ($\%$)Qwen 2.5
@@
Figure 4 plots mean RULER score (across all tasks) versus KV retention for Compactor and several baselines (SnapKV, TOVA, KeyDiff, PyramidKV, DuoAttention, random eviction) for both Llama 3.1 and Qwen 2.5 (no checkpoint for DuoAttention is available for Qwen 2.5). $100\%$ KV retention denotes the score of the model with no compression and is shown as a dashed line. We observe remarkably similar performance characteristics across both models. As the cache is pruned, Compactor degrades most gracefully: at $50\%$ retention it recovers $99\%$ of baseline, at $25\%$ it maintains $87.1\%$, and even at only $10\%$ retention it still achieves $68.0\%$. Performance when using SnapKV, TOVA, PyramidKV, and DuoAttention degrades poorly with KV retention setting, confirming the inadequacy of these methods in the query-agnostic regime. The performance of all methods is poor at severe compression ratios, demonstrating the inherent difficulty of the task. Performance of eviction methods on NIAH and QA tasks is shown in Figure 3, and demonstrates the variable performance across different tasks, which further motivates the need for context-calibrated compression.
@@
Single-Doc QA Compactor 44.6 44.5 41.9 35.8 30.0 42.0 41.6 38.3 33.3 25.3
KeyDiff 44.3 39.5 30.2 22.2 41.3 36.2 25.7 19.3
SnapKV 40.6 33.4 25.3 18.9 37.4 29.8 22.5 17.7
TOVA 41.8 32.7 22.1 15.8 38.4 29.9 19.2 15.4
Multi-Doc QA Compactor 39.3 38.8 36.4 32.8 27.9 46.1 45.1 45.4 39.4 33.5
KeyDiff 37.8 33.7 25.2 20.1 43.1 38.0 28.4 22.5
SnapKV 36.7 32.53 24.8 20.9 42.8 38.5 29.4 24.5
TOVA 36.7 34.2 25.6 20.6 44.5 41.4 31.7 24.5
Summarization Compactor 28.7 28.2 26.9 25.0 23.3 27.8 27.5 26.9 25.7 23.7
KeyDiff 28.2 26.7 23.4 21.3 27.2 25.3 22.1 19.9
SnapKV 27.1 24.9 22.5 20.2 26.5 24.6 22.3 20.0
TOVA 27.0 25.0 22.0 20.2 26.6 24.7 21.8 19.6
Few Shot Compactor 48.3 58.9 54.8 50.0 42.2 65.6 64.4 63.1 51.3 45.4
KeyDiff 56.3 54.9 49.9 44.1 64.4 61.5 53.8 46.5
SnapKV 50.8 53.0 48.4 45.1 64.4 62.9 56.4 50.8
TOVA 44.1 42.2 40.0 40.3 64.2 58.9 51.0 46.7
Code Compactor 43.1 45.1 44.4 44.0 42.9 62.4 62.2 61.2 55.6 53.2
KeyDiff 44.0 37.4 32.0 29.7 58.73 46.3 37.1 36.2
SnapKV 41.4 41.9 43.8 46.2 62.1 61.3 60.0 59.1
TOVA 43.4 43.4 44.0 45.8 62.2 61.9 59.4 58.3
Total Compactor 42.4 44.9 42.5 38.8 33.9 50.1 49.4 48.1 41.8 36.6
KeyDiff 43.9 40.0 33.1 27.7 48.3 42.9 34.2 29.4
SnapKV 41.2 38.5 33.3 29.7 47.9 44.1 38.0 33.6
TOVA 40.2 36.9 30.9 28.1 48.4 44.3 36.6 31.9
5.3 Longbench
We evaluate Compactor on Longbench to assess performance on real-world long context tasks. Table 1 reports per-task and average scores on LongBench for Compactor , TOVA, KeyDiff, and SnapKV at $50\%$, $25\%$, $10\%$, and $5\%$ retention. PyramidKV was excluded due to how similarly it performs to SnapKV. At $50\%$ retention, Compactor matches or slightly exceeds the full-cache performance, with particular gains in few-shot and coding tasks with less contextual diversity. Even at $25\%$, Compactor retains around $95\%$ of the full-cache score. Under more aggressive pruning ($10\%$ and $5\%$), Compactor ’s advantage grows, outperforming competing methods by over 15%. Furthermore, Compactor performs well for all task categories, whereas some methods greatly underperform (e.g KeyDiff on Code tasks, or TOVA on Single-Doc QA tasks). Such task-agnostic performance is essential in practical deployments of KV cache compression. These results confirm that Compactor reliably preserves tokens that drive task performance.
@@ 5.4 Context-Calibrated Compression
0202040406060808010010000.50.511UUID taskKV Retention ($\%$)NLL RatioLlama 3.1CompactorSnapKV0202040406060808010010000.50.511UUID taskKV Retention ($\%$)Qwen 2.5
Figure 5: NLL vs KV retention on 13 tasks from RULER, one line per sub-task. For each sub-task we compute the (reciprocal) gain in NLL of the ground-truth answer when conditioned on compressed contexts (for each compression method). This empirical trend motivates the exponential calibration function introduced in §3.5. These figures are generated without head-adaptive compression.
02020404060608080100100FinetunedZero ShotFull (100%)Full (100%) Compactor  (32%) SnapKV (51%) KeyDiff (48%) Random (90%) TOVA (53%) Compactor (19%)KeyDiff (32%)SnapKV (40%)Random (76%)TOVA (42%)KV Retention (%)
Figure 6: The top bar shows the KV retention rates that each compression method induces when using context-calibrated compression on Longbench. The bottom bar shows the same when the LLM is finetuned on documents from the Longbench test set (no queries). In all cases, the performance of the compression methods is within 0.1 of full KV cache performance ($42.4\pm 0.1$). Results are shown for Llama-3.1 8B.
@@
For all tested methods, the performance when applying context-calibrated compression is within $5\%$ of uncompressed cache performance. This confirms both the utility of NLL as a proxy for downstream task performance and the efficacy of the proposed calibration function at ensuring compression does not degrade performance. Note Compactor is able to achieve this performance while retaining fewer tokens. In particular, SnapKV’s is far removed from the excellent performance it demonstrates in the query-aware setting (li_snapkv_2024), where it can match baseline performance while retaining only $10\%$ of the cache.
Lastly, we finetune the model on documents from Longbench (no queries or answers), and evaluate the performance of context-calibrated compression with the finetuned model on Longbench (we evaluate and finetune with the same set of documents). We would expect that the NLL of the documents to fall, allowing for higher compression rates for the documents. Such a procedure is useful when we know a priori which corpus will be assessed, and we desire the highest possible accuracy and compression on contexts from this corpus. For example, a question-answering LLM may be finetuned on a corpus of supporting documents, while simultaneously being presented with relevant retrieved documents from that corpus at test time. The result of this experiment is shown in Fig 6 (“Finetuned"). As expected, we see that all models achieve higher compression rates when finetuned on the test corpus. Compactor still significantly outperforms competing techniques, though the gap is shortened; notably, Compactor without training (Zero Shot) outperforms SnapKV and TOVA even after finetuning